OpenStudy (anonymous):

Really need help! Been trying to solve for a while. 1. Solve the system by using a table. 3y=4x+7 , -4x-4y=28. These are the possible solutions:(–4, –3),(–3, –4), (4, 3),(3, 4). 2.Solve by graphing. -3x-y=-10, 4x-4y=8. possible solutions:(-1,3), (3-1), (1,3), (3,1)

OpenStudy (anonymous):

Also help with this problem : The spreadsheet shows the monthly revenue and expenses for a new business. Use the models to predict the month in which revenue will equal expenses. Possible answers: October September August May

OpenStudy (anonymous):

help with any of them would be really appreciated (:

OpenStudy (anonymous):

I think September is your answer for the second problem

OpenStudy (anonymous):

your estimate is as good as any haha. I have no clue so even an educated guess is better than nothing

OpenStudy (anonymous):

I honestly just suck horribly at solving systems of any sort x.x

OpenStudy (anonymous):

I answered it that way by reproducing the pattern of increase in the revenue and expense columns--that is, by adding differences in a row. Does that sound legitimate?

OpenStudy (anonymous):

Sounds alright to me haha

OpenStudy (anonymous):

I think I can solve the system, but I'm not sure what is meant by table. Is it all right if you use substitution and elimination?

OpenStudy (anonymous):

I would just really like some help with the 1st two because there are lot's of other questions like them, including on upcoming tests. That I really don't want to get wrong :/

OpenStudy (anonymous):

Yeah that's what I was trying to do. I however looked up solving solutions on google and it showed a way to do it where you used a graph to figure out where the two lines crossed. It didn't make sense with my equation though.

OpenStudy (anonymous):

\[-3x-y=-10\]\[ 4x-4y=8\] first equation is the same as \(y=-3x+10\) so you can replace \(y\) in the second equation by \(-3x+10\) with judicious use of parentheses you get \[4x-4(-3x+10)=8\] and now you have one equation and one variable to solve

OpenStudy (anonymous):

you get \[4x+12x-40=8\] \[16x-40=8\] \[16x=48\] \[x=48\div 16=3\]

OpenStudy (anonymous):

As long as it end up with an x and a y and they are one of the options above I really don't understand why it matters so much how we got it haha. I suppose in some cases it really matters but.. Thanks satellite that makes sense. I'll try writing it out.

OpenStudy (anonymous):

hope the steps are claere

OpenStudy (anonymous):

yeah I think so, so what next?

OpenStudy (anonymous):

now that you have \(x=3\) you can replace \(x\) by 3 in either equation to solve for \(y\)

OpenStudy (anonymous):

for example we know \(y=-3x+10\) and we know \(x=3\) making \(y=-3\times 3+10=-9+10=1\)

OpenStudy (anonymous):

okay haha, I think I get that too x.x

OpenStudy (anonymous):

solving by graphing is too much work, and besides, if the answer is \((-\frac{1}{7}, \frac{5}{7})\) for example you will never know it from the graph

OpenStudy (anonymous):

I know I really hate it too :/ Then again I hate all math. But it's what they are making us do so .

OpenStudy (anonymous):

yes, sometimes math (especially this kind) is hateful that is why they invented computers for example, here is the answer to the first one http://www.wolframalpha.com/input/?i=+3y%3D4x%2B7+%2C+-4x-4y%3D28

OpenStudy (anonymous):

oh wow that's really cool! Thanks alot.